# A Lightweight Approach for State Machine Replication

### v1

@@ Abstract
We present a lightweight solution for state machine replication with commitment certificates. Specifically, we adapt a simple median rule from the stabilizing consensus problem [19] to operate in a client-server setting where arbitrary servers may be blocked adaptively based on past system information. We further extend our protocol by compressing information about committed commands, thus keeping the protocol lightweight, while still enabling clients to easily prove that their commands have indeed been committed on the shared state. Our approach guarantees liveness as long as at most a constant fraction of servers are blocked, ensures safety under any number of blocked servers, and supports fast recovery from massive blocking attacks. In addition to offering near-optimal performance in several respects, our method is fully decentralized, unlike other near-optimal solutions that rely on leaders. In particular, our solution is robust against adversaries that target key servers (which captures insider-based denial-of-service attacks), whereas leader-based approaches fail under such a blocking model.
@@ 1 Introduction
In the state machine replication (SMR) problem [38], a set of servers is required to replicate a state machine in a consistent way by agreeing on an order in which commands from clients are to be committed on it. Formally, a state machine $(\mathcal{S},\mathcal{C})$ consists of a state space $\mathcal{S}$ and a collection of commands $\mathcal{C}$ acting on that state space. Each server maintains a copy of the state machine (simply called a shared state in the following). The clients may send their commands to any of the servers. The goal is to satisfy the following conditions:
@@
Safety and liveness imply that if every server executes the commands in the order in which they were committed by that server, the sequence of transitions on the shared state will be the same for each correct server, and any transition at one correct server will eventually be performed by all other correct servers. An important application of SMR is a system that manages financial transactions. However, for financial transactions it is not just important to maintain a shared state (i.e., the account balances) but also records of the transactions. In cryptocurrencies, transactions are typically recorded on a large blockchain that is replicated across many servers, for example, Bitcoin’s ledger exceeds 680 GB and is maintained by thousands of full nodes [27]. A much more space-efficient alternative for servers would be to maintain only the shared state and shift the burden of maintaining a proof that a transaction took place to the originating clients, who are naturally incentivized to do so, to be able to prove successful commitment of their transactions to third parties.
@@
In this paper, we present a scalable and robust solution to the SMR problem that provides safety and liveness under a blocking adversary and minimizes the amount of information that must be maintained by the servers by offloading the responsibility of certifying the commitment of commands to the clients. Specifically, we require clients to certify commitment of their commands to other clients, without using long-term cryptographic signatures, where servers may act only as verifiers that retain a small amount of metadata but not the complete history of committed commands. This poses several intertwined challenges, especially under failures. We adopt a blocking-failure model under an adaptive adversary that has full information with a one-round delay. Adaptive blocking failures subsume both crash-stop and crash–recovery behavior since such an adversary may permanently or temporarily block servers. The core difficulty is that the adversary can potentially isolate groups of servers from each other by alternate blockings, which can make it hard to maintain safety.
Unlike many contemporary solutions for fault-tolerant SMR, which use leaders that can be easily suppressed by our adaptive adversary, our protocol is leaderless and, to our knowledge, the first to realize SMR via the median-rule consensus dynamics [19]. We rigorously prove safety and liveness properties of our protocol under scenarios with increasing difficulty, for which we provide a formal framework by defining a sequence of core problems building on top of each other. In Section 1.2 we start with the basic consensus problem with limited server blockings and extend it, step by step, to the most challenging scenario of SMR with lightweight servers under unlimited server blockings. Our final protocol guarantees liveness with small bounds on latency and bandwidth per command (logarithmic in $n$), supports signature-free commitment certificates with lightweight servers, and provides safe recovery even after an arbitrary number of server blockings (up to all). Along the way, we develop new analysis tools for median-like rules under blocking adversaries, identifying phenomena (e.g., “death spirals”) that we must avoid or even carefully exploit. Our solution is randomized, but its failure probability can be tuned to be vanishingly small, rendering violations practically negligible. This work focuses on the theoretical foundations of light-weight SMR under a blocking adversary based on the median rule, but since the underlying algorithms are elementary, we expect them to be useful in practice.
@@ 1.1 The Model
Communication between clients and servers is done by point-to-point message passing. While clients can be addressed directly, for servers we assume the weaker gossip communication model (see, e.g., [26]) where servers can exchange a point-to-point message with anyone but cannot choose whom to contact. Messages are exchanged using either a push operation, which sends a message to a random server, or a pull operation, which asks a random server to reply to the sender, where every server has the same probability to be chosen as the target of the operation. Gossip-based protocols are highly practical due to their inherent load balancing, rapid convergence, simplicity, and robustness under stress or disruption. For instance, they play a key role in Avalanche consensus [3, 37] and anonymous networks like Tor. Because the gossip model is used for the servers, neither servers nor clients need to know the server identifiers or the exact number $n$; having some polynomial estimate of $n$ is sufficient for our protocols to work.
@@
We assume that the clients cannot impersonate other clients but may otherwise behave in a Byzantine manner. Servers are subject to blocking failures that are adversarially triggered but correctly follow the given protocol. The blocking adversary can isolate any server at any point in time, preventing it from sending or receiving messages in a given round. A server does not know whether it is isolated or not, and in case of a failed push or pull request, no information is given who of the two involved parties caused the failure. This covers real-world phenomena such as denial-of-service attacks or network failures. It also captures aspects of network asynchrony or unreliable links, in the sense that blocked parts of the network experience delays or message loss. Because blocked servers are effectively cut out of the protocol, this model also subsumes the crash-recovery model, with the adversary controlling the timing of crashes and subsequent recovery. We consider a late adversary that adapts its blocking decisions based on previously observed system states (e.g. [2, 36]) which allows us to rely on randomness to achieve network-wide outcomes that are outside the adversary’s control. Lateness is a natural assumption, as information about the servers first has to be aggregated and evaluated by an adversary in order to target its attack.
@@
We assume a 1-late blocking adversary ($\alpha=1$): in round $t$ it knows the entire system state only up to the beginning of round t1t-1. This model is incomparable to, and not weaker than, the classical Byzantine model. Byzantine-tolerant protocols typically posit a fixed, sufficiently large correct majority that is always available, whereas here the adversary may adaptively block any subset of servers at any time. In particular, leader-based consensus protocols cannot maintain liveness against such adaptive blocking.
More generally, for an $\alpha$-late adversary with $\alpha\geq 1$, the adversary cannot observe which client requests arrive in the current round, nor the recipients of the servers’ push/pull messages from the previous round, and therefore, it cannot reliably infer the servers’ current states. This one-round opacity is crucial for fast commitment, because with $\Theta(n)$ adaptive crash failures under a full-information adversary, any synchronous randomized consensus protocol needs $\tilde{\Theta}(\!\sqrt{n})$ rounds [9].
@@
We say an event occurs with high probability (w.h.p.) when the probability of failure is at most $1/n^{c}$ for any constant $c>0$. We aim at algorithms that succeed w.h.p., ensuring that as $n$ becomes sufficiently large, the algorithm satisfies our safety and liveness requirements within any reasonable amount of time. Further probabilistic concepts that we use frequently are listed in Appendix A.
@@ 1.2 Overview of our Approach
Note that a protocol ensuring agreement should eliminate values from blocked servers to prevent the survival of outdated ones. Our agreement requirement for useful servers is analogous to Byzantine agreement (BA), where only correct servers must agree. The key difference is that the set of useful servers changes over time, due to the adaptive adversary and the randomness of gossip. A main challenge, therefore, is to ensure that a constant fraction of servers remains useful (available) at all times despite this shifting membership. We address this problem in Section 2 by combining a simple median rule in [19] with undecided-state dynamics (see, e.g., [4]). In this median rule, each server in each round requests the value from a random, constant-size subset of other servers, adopts the median value if enough values are received, and otherwise deletes its value, thereby becoming useless (i.e., not useful).
@@
The clear drawback of deleting committed commands is that it erazes the history of commands (e.g., transactions in a blockchain). This is the next problem we address: preserving the history without the servers needing to store it. Our solution, given in Section 5, is to offload the main burden of storing committed commands to the clients that issued them together with a small certificate that allows clients to prove to any third party that their command is indeed part of the committed history, whereas the servers only store a small amount of metadata to verify certificates. Note that our approach makes the client completely responsible for storing its issued commands and the associated certificate. However, loss of this data by one client never affects certificates of others.
@@
Technically, we combine the median rule of [19] with undecided-state dynamics (e.g., [4]) and extend it to SMR with lightweight servers. To our knowledge, this is the first use of median-rule dynamics for SMR and the first analysis under a late blocking adversary. Prior work largely targets Byzantine faults and has either higher overheads or is leader-based and thus cannot ensure liveness against our adaptive blocking adversary.
@@
A further contribution is that servers do not store the full command history, as is common in blockchains. Instead, commitment certificates are offloaded to the issuing clients. Servers keep only the shared state and the pool of uncommitted commands. Clients need certificate updates only when they submit a new command, and can remain offline otherwise. Our approach has additional benefits to privacy: after a command is committed, only the client will retain command-specific information. Technically, the recovery protocol is the most delicate component. Our techniques require new insights into the median rule under blockings and two tightly coupled subprotocols, enabling rapid (logarithmic-time) recovery after arbitrary blocking attacks.
@@ 1.3 Related Work
Consensus Protocols and Modeling Assumptions
A common approach to state-machine replication (SMR) is to use a Byzantine fault-tolerant (BFT) consensus protocol. The formal study of BFT protocols dates back to Lamport’s pioneering work [34] and has since been the focus of extensive research, as covered in various textbooks (e.g., [7, 13]). Historically, BFT protocols saw limited adoption since the overhead was considered too high. The emergence of blockchains, however, triggered new efforts to make these protocols practical. Blockchain-oriented consensus protocols are typically categorized into permissioned and permissionless models. Prominent permissionless protocols include proof of work [33] and proof of stake [22, 30]. Although they are central to many cryptocurrencies like Bitcoin and Ethereum, we do not discuss them here, as our focus lies on permissioned systems.
BFT protocols for permissioned blockchains assume a fixed set of $n$ participants who know each other, with an adversary capable of controlling up to $f$ of them in any manner (see [44] for a recent survey). Since even a single fault can undermine safety and liveness in the asynchronous model [21], the gold standard is to guarantee liveness with small latency in a partially synchronous environment [20], while preserving safety under asynchrony. Partial synchrony has two flavors: synchrony after an unknown global stabilization time, or delays bounded by an unknown $\Delta$. Under reasonable assumptions, a protocol robust to one can be adapted to the other [16].
Comparison of Leader and DAG Based Consensus Protocols
@@
Most practical BFT consensus protocols follow either a leader-based or a DAG-based approach. Examples of leader-based protocols are PBFT [14], HotStuff [43], Autobahn [24], and ProBFT [8], and examples of DAG-based protocols are DAG-Rider [29], Narwhal [18], Bullshark [39], and Shoal++ [5]. Most of the BFT protocols can tolerate large-scale Byzantine behavior in the network. However, with trusted components, such as those available today in trusted execution environments (TEEs), such strong fault tolerance might not be strictly necessary: adversarial peers can block a trusted device or obstruct message delivery, but cannot tamper with its internal execution. This has led to efficient consensus protocols based on trusted components [15, 42, 31], even before the blockchain era.
@@
All of the works above assume that the participants under adversary control is static. Although we consider only attacks that block participants, the blocking can change adaptively from round to round. This rules out leader-based approaches since such an adversary can simply block a leader once it has been elected with a denial-of-service (DoS) attack. Heavily concentrating communication on a single participant also leads to imbalances in communication load and raises fairness concerns. Although rotating the leader over time can address fairness, our protocol avoids leaders altogether by being fully decentralized, ensuring fairness in every round. In addition, while many state-of-the-art protocols rely on batching, coding, or threshold signatures to achieve near-optimal message complexity, our approach achieves near-optimal performance without such mechanisms.
@@
The primary advantage of DAG-based protocols is liveness under full asynchrony, whereas our protocol offers liveness only under partial synchrony, when enough participants are synchronous. However, DAG-based approaches can make very slow progress, a classical lower bound for randomized consensus in the asynchronous model shows that, for any integer $k$, an $f$-resilient algorithm over $n$ processes fails to terminate within k(nf)k(n-f) steps with probability at least $1/c^{k}$ (for some constant $c$) [6]. Thus, asynchrony must persist for a long time before one can benefit from any protocol that offers liveness under asynchrony, whereas our protocol recovers quickly once a period of arbitrary asynchrony ends.
Comparison of Gossip Based Consensus Protocols
@@
In contrast to leader-based or DAG-based protocols, our approach relies on the gossip communication model. Gossip-based protocols have been studied extensively for various tasks, including information dissemination [28], aggregation [26], network coding [25], and consensus [19]. See [10] for a survey on gossip-based protocols. Due to their fast convergence and resilience under stress and disruptions, they have also been employed in various blockchain solutions, such as Bitcoin [17] and Tendermint [12], for network maintenance and information dissemination. Most closely aligned with our work are anti-entropy protocols [11, 35, 41], including the Snow consensus protocol used in Avalanche [37]. While many of these protocols have been analyzed under message loss and some under Byzantine behavior (e.g., [3, 37]), their efficiency and robustness are often evaluated primarily through simulations. In contrast, we provide a rigorous analysis, building on works such as [19, 36]. For the gossip model, it is known that under $\Theta(n)$ adaptive crash failures by a full-information adversary, any synchronous randomized consensus protocol needs $\tilde{\Theta}(\sqrt{n})$ rounds [9]. We bypass this lower bound by withholding new information from the adversary for one round.
@@
Our approach to generate client certificates is based on a Merkle hash forest, which builds on the Merkle Mountain Range technique used in Ethereum and other blockchains [40]. While methods to store a compressed representation of the Merkle hash forest on servers have been used before, the novel challenge we address is ensuring that clients only need to receive the necessary hashing information once, at the moment a command is acknowledged by a server, yet are still able to prove to any useful server, at any later time, that their command was committed, even though the servers just store a compressed version of the Merkle hash forest.

### v2

@@ Abstract
We present a lightweight solution for state machine replication with commitment certificates. Specifically, we adapt and analyze a median rule for the stabilizing consensus problem [DoerrGMSS11] to operate in a client-server setting where arbitrary servers may be blocked adaptively based on past system information. We further extend our protocol by compressing information about committed commands, thus keeping the protocol lightweight, while still enabling clients to easily prove that their commands have indeed been committed on the shared state. Our approach guarantees liveness as long as at most a constant fraction of servers are blocked, ensures safety under any number of blocked servers, and supports fast recovery even after all servers are blocked. In addition to offering near-optimal asymptotic performance in several respects, our method is fully decentralized, unlike other near-optimal solutions that rely on leaders. In particular, our solution is robust against adversaries that target key servers (which captures insider-based denial-of-service attacks), whereas leader-based approaches fail under such a blocking model.
Keywords: analysis of consensus dynamics, distributed algorithms, state machine replication
@@ 1 Introduction
In the state machine replication (SMR) problem [Schneider90], a set of servers is required to replicate a state machine in a consistent way by agreeing on an order in which commands from clients are to be committed on it. Formally, a state machine $(\mathcal{S},\mathcal{C})$ consists of a state space $\mathcal{S}$ and a collection of commands $\mathcal{C}$ acting on that state space. Each server maintains a copy of the state machine (simply called a shared state in the following). The clients may send their commands to any of the servers. The goal is to satisfy the following conditions:
@@
Safety and liveness imply that if every server executes the commands in the order in which they were committed by that server, the sequence of transitions on the shared state will be the same for each correct server, and any transition at one correct server will eventually be performed by all other correct servers. An important application of SMR is a system that manages financial transactions. However, for financial transactions, it is not only important to maintain a shared state (i.e., the account balances), but also records of the transactions. In cryptocurrencies, transactions are typically recorded on a large blockchain that is replicated across many servers. For example, Bitcoin’s ledger exceeds 699 GB and is maintained by thousands of full nodes [BBS]. A much more space-efficient alternative for servers would be to maintain only the shared state and shift the burden of maintaining a proof that a transaction took place to the originating clients, who are naturally incentivized to do so, to be able to prove successful commitment of their transactions to third parties.
@@
We present a scalable and robust solution to the SMR problem that provides safety and liveness under a blocking adversary and minimizes the amount of information that must be maintained by the servers by offloading the responsibility of certifying the commitment of commands to the clients. Specifically, we require clients to certify commitment of their commands to other clients, without using long-term cryptographic signatures, where servers may act only as verifiers that retain a small amount of metadata but not the complete history of committed commands. This poses several intertwined challenges, especially under failures. We adopt a blocking-failure model under an adaptive adversary that has full information with a one-round delay. Adaptive blocking failures subsume both crash-stop and crash–recovery behavior since such an adversary may permanently or temporarily block servers. The core difficulty is that the adversary can potentially isolate groups of servers from each other by alternate blockings, making it hard to maintain safety.
Unlike many contemporary solutions for fault-tolerant SMR, which use leaders that can be easily suppressed by our adaptive adversary, our protocol is leaderless and, to our knowledge, the first to realize SMR via the median-rule consensus dynamics [DoerrGMSS11]. We prove safety and liveness properties of our protocol in scenarios with increasing difficulty, for which we provide a formal framework by defining a sequence of core problems built on top of each other. This work focuses on the theoretical foundations of SMR under a blocking adversary based on the median rule by providing an analysis of the consensus dynamics in the style of [BecchettiCN20, DoerrGMSS11] and using these insights to obtain a light-weight protocol that is robust to blocking failures.
@@ 1.1 The Model
Communication between clients and servers is done by point-to-point message passing. While clients can be addressed directly, for servers we assume the weaker gossip communication model (see, e.g., [DBLP:conf/podc/HaeuplerMS18]) where servers can exchange a point-to-point message with anyone but cannot choose whom to contact. Messages are exchanged using either a push operation, which sends a message to a random server, or a pull operation, which asks a random server to reply to the sender, where every server has the same probability to be chosen as the target of the operation. Gossip-based protocols are highly practical due to their inherent load balancing, rapid convergence, simplicity, and robustness under stress or disruption. For instance, they play a key role in Avalanche consensus [AmoresSesarCS24, abs-1906-08936] and anonymous networks like Tor. Because the gossip model is used for the servers, neither servers nor clients need to know the server identifiers or the exact number $n$; having some polynomial estimate of $n$ is sufficient for our protocols to work.
@@
We assume that the clients cannot impersonate other clients but may otherwise behave in a Byzantine manner. Servers are subject to blocking failures that are adversarially triggered but otherwise correctly follow the given protocol. The blocking adversary can isolate any server at any point in time, preventing it from sending or receiving messages in a given round. A server does not know whether it is isolated or not, and in case of a failed push or pull request, no information is given on who of the two involved parties caused the failure. This covers real-world phenomena such as denial-of-service attacks or network failures. It also captures aspects of network asynchrony or unreliable links, in the sense that blocked parts of the network experience delays or message loss. Because blocked servers are effectively cut out of the protocol, this model also subsumes the crash-recovery model, with the adversary controlling the timing of crashes and subsequent recovery. We consider a late adversary that adapts its blocking decisions based on previously observed system states (e.g. [DBLP:conf/opodis/AhmadiGKM15, RobinsonSS18]) which allows us to rely on randomness to achieve network-wide outcomes that are outside the adversary’s control. Lateness is a natural assumption, as information about the servers first has to be aggregated and evaluated by an adversary in order to target its attack.
@@
We assume a 1-late blocking adversary ($\alpha=1$): in round $t$ it knows the entire system state only up to the beginning of round t1t-1. This model is incomparable to, and not weaker than, the classical Byzantine model. Byzantine-tolerant protocols typically posit a fixed, sufficiently large correct majority that is always available, whereas here the adversary may adaptively block any subset of servers at any time. In particular, leader-based consensus protocols cannot maintain liveness against such adaptive blocking. More generally, for an $\alpha$-late adversary with $\alpha\geq 1$, the adversary cannot observe which client requests arrive in the current round, nor the recipients of the servers’ push/pull messages from the previous round, and therefore, it cannot reliably infer the servers’ current states. This one-round opacity is crucial for fast commitment, because with $\Theta(n)$ adaptive crash failures under a full-information adversary, any synchronous randomized consensus protocol needs $\tilde{\Theta}(\!\sqrt{n})$ rounds [Bar-JosephB98].
@@
We use the standard concept of with high probability (w.h.p.) where the total probability of algorithm failure is upper bounded by $1/n^{c}$ for any constant $c>0$. This ensures that the likelihood that our algorithm does not satisfy the safety and liveness vanishes for sufficiently large $n$. We further discuss how to deal with the unlikely event of a single failure in Section 7. The probabilistic concepts we frequently use are given in Appendix A.
@@ 1.2 Overview of our Approach
Note that a protocol ensuring agreement should eliminate values from blocked servers to prevent the survival of outdated ones. Our agreement requirement for useful servers is analogous to Byzantine agreement (BA), where only correct servers must agree. The key difference is that the set of useful servers changes over time, due to the adaptive adversary and the randomness of gossip. A main challenge, therefore, is to ensure that a constant fraction of servers remains useful (available) at all times despite this shifting membership. We address this problem in Section 2 by combining a simple median rule in [DoerrGMSS11] with undecided-state dynamics (see, e.g., [AngluinAE08]). In this median rule, each server in each round requests the value from a random, constant-size subset of other servers, adopts the median value if enough values are received, and otherwise deletes its value, thereby becoming useless (i.e., not useful).
@@
The clear drawback of deleting committed commands is that it erases the history of commands (e.g., transactions in a blockchain). This is the next problem we address: preserving the history without the servers needing to store it. Our solution, given in Section 5, is to offload the main burden of storing committed commands to the clients that issued them together with a small certificate that allows clients to prove to any third party that their command is indeed part of the committed history, whereas the servers only store a small amount of metadata to verify certificates. Note that our approach makes the client completely responsible for storing its issued commands and the associated certificate. However, loss of this data by one client never affects certificates of others.
@@
Technically, we combine the median rule of [DoerrGMSS11] with undecided-state dynamics (e.g., [AngluinAE08]) and extend it to SMR with lightweight servers. To our knowledge, this is the first use of median-rule dynamics for SMR and the first analysis under a late blocking adversary. Prior work largely targets Byzantine faults and has either higher overheads or is leader-based and thus cannot ensure liveness against our adaptive blocking adversary.
@@
Further, we show that servers do not need to store the full command history, as is common in blockchains. Instead, commitment certificates are offloaded to the issuing clients. Servers keep only the shared state and the pool of uncommitted commands. Clients need certificate updates only when they submit a new command, and can remain offline otherwise. Our approach has additional benefits to privacy: after a command is committed, only the client will retain command-specific information. Technically, the recovery protocol is the most delicate component. Our techniques require new insights into the median rule under blockings and two tightly coupled subprotocols, enabling rapid (logarithmic-time) recovery after arbitrary blocking attacks.
@@ 1.3 Related Work
Leader and DAG Based Consensus Protocols
@@
Most practical Byzantine fault-tolerant (BFT) consensus protocols follow either a leader-based or a DAG-based approach. Examples of leader-based protocols are PBFT [DBLP:conf/osdi/CastroL99/PBFT], HotStuff [YinMRGA19], Autobahn [DBLP:conf/sosp/GiridharanSAAC24], and ProBFT [DBLP:conf/podc/AvelasHADB24], and examples of DAG-based protocols are DAG-Rider [DBLP:conf/podc/KeidarKNS21], Narwhal [DBLP:conf/eurosys/DanezisKSS22], Bullshark [SpiegelmanGSK22], and Shoal++ [Shoal++]. Most of the BFT protocols can tolerate large-scale Byzantine behavior in the network. However, with the advent of trusted execution environments (TEEs), such strong fault tolerance might not be strictly necessary: adversarial peers can block a trusted device or obstruct message delivery, but cannot tamper with its internal execution. This has led to efficient consensus protocols based on trusted components [DBLP:conf/sosp/ChunMSK07, DBLP:conf/nsdi/LevinDLM09, DBLP:journals/tc/VeroneseCBLV13], even before the blockchain era.
@@
All of the works above assume that the set of participants under adversary control is static. Although we consider only attacks that block participants, the blocking can change adaptively from round to round. This rules out leader-based approaches since such an adversary can simply block a leader once it has been elected with a denial-of-service (DoS) attack. Heavily concentrating communication on a single participant also leads to imbalances in communication load and raises fairness concerns. Although rotating the leader over time can address fairness, our protocol avoids leaders altogether by being fully decentralized. In addition, while many state-of-the-art protocols rely on batching, coding, or threshold signatures to achieve near-optimal message complexity, our approach achieves near-optimal performance without such mechanisms.
@@
The primary advantage of DAG-based protocols is liveness under full asynchrony, whereas our protocol offers liveness only under partial synchrony, when enough participants are synchronous. However, DAG-based approaches might make very slow progress: a classical lower bound for randomized consensus in the asynchronous model shows that, for any integer $k$, an $f$-resilient algorithm over $n$ processes fails to terminate within k(nf)k(n-f) steps with probability at least $1/c^{k}$ (for some constant $c$) [AttiyaH10]. Thus, asynchrony must persist for a long time before one can benefit from any protocol that offers liveness under asynchrony, whereas our protocol recovers quickly once a period of arbitrary asynchrony ends.
Gossip Based Consensus Protocols
@@
In contrast to leader-based or DAG-based protocols, our approach relies on the gossip communication model. Gossip-based protocols have been studied extensively for various tasks, including information dissemination [DBLP:conf/focs/KarpSSV00], aggregation [DBLP:conf/podc/HaeuplerMS18], network coding [DBLP:conf/stoc/Haeupler11], and consensus [DoerrGMSS11]. See [BecchettiCN20] for a survey on gossip-based protocols. Due to their fast convergence and resilience under stress and disruptions, they have also been employed in various blockchain solutions, such as Bitcoin [DBLP:conf/icdcn/CrucianiP23] and Tendermint [DBLP:journals/corr/abs-1807-04938], for network maintenance and information dissemination. Most closely aligned with our work are anti-entropy protocols [BirmanHOXBM99, PetersenSTTD97, RenesseDGT08], including the Snow consensus protocol used in Avalanche [abs-1906-08936]. While many of these protocols have been analyzed under message loss and some under Byzantine behavior (e.g., [AmoresSesarCS24, abs-1906-08936]), their efficiency and robustness are often evaluated primarily through simulations. In contrast, we provide a rigorous analysis, building on works such as [DoerrGMSS11, RobinsonSS18]. For the gossip model, it is known that under $\Theta(n)$ adaptive crash failures by a full-information adversary, any synchronous randomized consensus protocol needs $\tilde{\Theta}(\sqrt{n})$ rounds [Bar-JosephB98]. We bypass this lower bound by withholding new information from the adversary for one round.
@@
Our approach to generate client certificates is based on a Merkle hash forest, which builds on the Merkle Mountain Range technique used in Ethereum and other blockchains [PeterTodd12]. While methods to store a compressed representation of the Merkle hash forest on servers have been used before, the novel challenge we address is ensuring that clients only need to receive the necessary hashing information once, at the moment the commitment of a command is acknowledged by a server, yet are still able to prove to any useful server, at any later time, that their command was committed, even though the servers just store a compressed version of the Merkle hash forest.